Fermi National Laboratory
 

Accelerator Update for March 23, 2007 - March 26, 2007

 

Friday March 23

The day shift began with Operations monitoring store 5296 (87.51E30) the Recycler stash (181.9E10), with the Antiproton Source (Pbar) stacking (59.7E10), with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

Operations established beam to the Meson MTest experiment at 8:25 AM.

At 8:33 AM, Operations lowered the Booster beam intensity because of the many MI beam permit trips caused by bad beam positions.

Operations prepared Booster for access at 9:28 AM.

Operations prepared the Meson M01 area for access at 10:11 AM.

A Linac expert began a study on a RF station (LRF3) at 11:40 AM. The study took an hour.

Booster was out of access by 12:33 PM.

Pbar, NuMI, and MTest resumed taking beam at 1:08 PM.

At 3:18 PM, the MIRF system tripped off due to a high temperature alarm on the 95 LCW system. An expert investigated. He found that the LCW system was all right, but that the temperature sensor had a bad connector, which he fixed.

Operations reported at 4:16 PM that the Recycler had stashed antiprotons once during the day shift. Recycler transverse cooling problems held off stashing for most of the shift. Experts were investigating.

The MTest experiment finished using the beam for the night at 6:08 PM.

Operations reported at 11:52 PM that the Recycler had stashed four times during the evening shift.

Saturday March 24

The midnight shift began with Operations monitoring store 5296 and the stash (345.4E10), with Pbar stacking (15.4E10), and with NuMI taking beam.

Operations terminated store 5296 at 12:03 AM.

Operations put the Tevatron (TeV) into shot setup at 12:14 AM.

Operations established store 5297 at 2:18 AM with an initial luminosity of 242.2E30.

At 4:10 AM, the TeV sectors B3 air compressor tripped off. Operators reset the compressor, but noted its motor was very hot. Operations informed experts.

At 5:19 AM, Operations reported that the MI beam permit had been tripping off due to bad beam positions.

At 6 AM, BRF5 tripped off. It reset.

Operations established beam to MTest at 6 AM.

At 6:51 AM, the Linac 750keV area buncher tripped off with no bad status. It reset, but would not run up remotely. Operators ran it up locally.

Operations reported at 8:29 AM that the Recycler had stashed four times during the midnight shift.

At 8:55 AM, the Booster West Anode Power Supply (WAPS) tripped off due to 750 KeV buncher trip. Everything reset. BRF10 also tripped, but would not reset. Operations bypassed the station.

The MTest experiment requested an 8 GeV tune at 8:59 AM. Operators tried to establish the beam, but couldn’t. Operations called in a Switchyard (SY) expert.

At 9:44 AM, Operations reported that the H- Source had been sparking since 8 AM, which had been causing Booster beam position problems.

At 9:48 AM, Operations reported that the SY expert was unable to establish beam to MTest. The investigation continues.

At 11:25 AM, Operations contacted Recycler experts about beam loss and high emittances. Operators tried some suggested changes with no help. An expert determined that the emittance blow up was due to cooling problems. Operators found some bad settings.

At 1:47 PM, Operations reported that the SY expert managed to get some beam to the MTest expert, but not the amount they wanted.

At 2:01 PM, LRF3 tripped off twice in 45 minutes. It reset.

At 3:03 PM, MI aborted the beam due to bad beam positions for the first of many, many times over four hours. It reset each time

At 4:39 PM, the Booster WAPS tripped off on an AC input. Operations had the power supply up in five minutes.

At 5:23 PM, BRF8 tripped off. Operators tried to reset the station, but due to sparking it wouldn’t stay on. Operators bypassed the station and reduced the beam intensity.

At 6:11 PM, Operations reported that the Recycler continued to have transverse emittance problems.

At 6:33 PM, operators tried to turn on BRF10, but this caused WAPS to trip off.

An RF expert arrived to investigate problems with BRF8 & 10 at 7:08 PM. He reported that BRF10 had a screen short on its power amplifier (PA) tube, which would require a tunnel access.

Operations reported at 7:59 PM that the Recycler had stashed four times during the last 12 hours. Store 5297 had a luminosity of 56.79E30, the stack was at 23.8E10, and the stash was at 210.3E10.

MTest finished using the beam at 8:03 PM.

At 8:27 PM, the TeV suffered a sectors E1 quench.

Operations prepared CDF and D0 for access at 9:35 PM.

MTest finished using the beam at 9:52 PM.

RF experts reported at 10:01 PM that the BRF8 needed its modulator replaced. The Crew Chief gave them permission to begin the repair.

At 10:09 PM, Operations reported that the MI beam permit tripped off for the first of six times in 45 minutes due to bad beam positions. It reset.

At 10:20 PM, Operations reported that BRF7 had tripped off a couple of times. It reset.

CDF came out of access at 10:21 PM.

A TeV expert reported at 10:28 PM that they found a vertical distortion to the TeV orbit prior to the quench. Operators replaced a card for a power supply (T:VA11).

D0 came out of access at 11:37 PM.

RF experts reported at 11:42 PM that they had completed the modulator change for BRF8. Booster beam resumed at 11:56 PM.

Sunday March 25

The midnight shift began with the TeV turn on in progress, with Operations monitoring the stash, with Pbar stacking, and with NuMI taking beam.

At 12:09 AM, MI tripped off due to bad beam position for the first of six times in six hours. It reset.

Operations loaded a 4x0 wet store into the TeV at 1:32 AM.

Operations put the TeV into shot setup at around 1:46 AM.

Operations established store 5299 at 4:35 AM with an initial luminosity of 209.6E30.

Recycler experts began an Electron Cooling (ECool)beam alignment at 4:38 AM.

Operations established beam to MTest at 6:19 AM.

At 6:39 AM, LRF3 tripped off. It reset.

At 8:20 AM, Operations reported that SY tripped off five times over the next nine hours. It reset.

Operations reported at 8:28 AM that the Recycler had stashed twice during the last 12 hours. Store 5299 had a luminosity of 116.7E30, the stack was at 110.8E10, and the stash was at 2.8E10. The odd numbers for the stack and stash were due to ECool beam alignment.

At 8:20 AM, Operations reported that SY tripped off five times over the next nine hours. It reset.

At 9:24 AM, the 750 KeV area buncher tripped off. Operators reset the buncher and beam resumed about five minutes later.

At 4:51 PM, Operations reported that the Booster Low-Level RF room temperature had gone into alarm. The Duty Electrician reported that the room’s thermostat had failed. They put a spot cooler in the room while they continued their investigation.

At 5:12 PM, Operations reported Booster and MI beam problems. After some investigation, the bad beam positions were traced to the LLRF room temperature problems.

MTest finished using the beam at 6:03 PM.

At 7:11 PM, LRF3 tripped off. It reset.

Operations reported at 8:19 PM that the Recycler had stashed five times during the last 12 hours. Store 5299 had a luminosity of 51.15E30, the stack was at 21.2E10, and the stash was at 208.2E10.

At 8:34 PM, Operations reported that the Pelletron had tripped off at the end of the last shift. An expert had tried to help operators get it back on, but it would trip back off. An expert was called.

The Duty Electrician reported at 9:01 PM that he had repaired the air conditioning and the Booster LLRF room temperature was under control.

At 9:44 PM, LRF3 tripped off for the first of four times in about 30 minutes.

At 10:51 PM, Operations reported that one or two batches were lost out of the last Recycler stash. Operators investigated and found that the Recycler cooling gating had turned itself on. Operators turned it off.

Monday March 26

The midnight shift began with Operations monitoring store 5299 and the stash, with Pbar stacking, and with NuMI taking beam.

At 12:39 AM, LRF3 tripped off. It reset.

At 12:43 AM, the Pelletron tripped off just before a stash. Operations contacted an expert. It reset.

At 2:10 AM, the NuMI horn tripped off. It reset.

At 2:24 AM, the Pelletron tripped off again. Operations contacted an expert who first helped operators get it back on and then came in to investigate further.

Operations terminated store 5299 at 3:59 AM.

Operations put the TeV into shot setup at 4:21 AM.

At 6:05 AM, a Pbar Debuncher RF station (DRF1-2) tripped off. It reset okay, but the cavity took about 30 minutes to recover its phase.

Operations established store 5300 at 6:19 AM with an initial luminosity of 235.7E30.

MTest resumed taking beam at 7:30 AM.

Operations reported at 8:26 AM that the Recycler had stashed four times during the last 12 hours.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Pbar experts to conduct studies this morning and for Recycler experts to conduct studies this afternoon.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/26/2007    email Fermilab